OpenStudy (anonymous):

The ratio of the sums of m and n terms of an A.P. is m^2: n^2. Show that the ratio of mth and nth term is (2m – 1): (2n – 1).

hartnn (hartnn):

u need to do some algebra here .... i assume you know the formulas foe sum and n'th term of AP did u reach to a point where u get 2a=d ??

hartnn (hartnn):

write formulas for sum , cancel m and n , cross-multiply, u get (2a-d)n=(2a-d)m means 2a=d now write the n'th term formual put d=2a and cancel a from numerstor and denom., u'll get you result.

hartnn (hartnn):

sorry for typos.....

OpenStudy (anonymous):

Well, that also means m=n .. And thus, ratio of two same things is 1 .. No ?

hartnn (hartnn):

No.

hartnn (hartnn):

if (2a-d)n =(2a-d)m since m and n are distinct, 2a-d = 0 will always satisfy (2a-d)n =(2a-d)m so, 2a=d

hartnn (hartnn):

u can look it this way also 2an-nd = 2am-md 2a(n-m)=d(n-m) so dividing by m-n as m-n\(\ne\)0 we get 2a=d

hartnn (hartnn):

still not convinced ?

OpenStudy (anonymous):

Sort of! :P

hartnn (hartnn):

did u get the ratio as (2m – 1): (2n – 1) and explicitly mention where u have doubt ?

OpenStudy (anonymous):

Nai nai nai, I AM convinced :D THANK YOU! :P :D

hartnn (hartnn):

welcome ^_^